Jan is 8 years older than Natalie. If the sum of their ages is 42, find their ages.
Answer by jndarrell(58)

Natalie = x years
Jan = x+8 years
Sum = addition
x + x+8 = 42
Simplify
2x+8=42
Subtract the 8 from both sides of the equation as the first step to get x alone
2x=34
Divide both sides by 2 to get the x alone
x=17
The question asked you for both ages, plug the solution to x in our set up at the top *DONT FORGET UNITS*
Natalie: 17years
Jan: 25 years